OpenStudy (anonymous):

A puppy weighed 27 lb. His weight increased by 233% by the time he was 1 year old. How much did the puppy weigh when he was 1 year old?

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (vianne):

1. you multiple 27 by 2.33 (which is the percent in the decimal form) 2. you add 27 to (27x2.33) 3. You should get about 89.91 pounds

OpenStudy (butterflydreamer):

hmmm. Well we know original he weighed 27 lb right? This means at 100% = 27 lb We know that there was a 223% INCREASE in the puppy's weight (daamn LOL) So that means we need to find out what 233 % + 100 % = 333 % equals to. To do this, first find out what 1% equals to. We know 100% = 27 lb. 1% = 27/100 = ? Then multiply this number by 333 to get 333% :)

OpenStudy (vianne):

Wait, are we trying to get percentage format or LB?

OpenStudy (butterflydreamer):

lb because the question asks for the weight of the puppy :)

OpenStudy (anonymous):

0.27

OpenStudy (vianne):

Ahh right.

OpenStudy (butterflydreamer):

yeah so 1% = 0.27 333% = 0.27 * 333 = ?

OpenStudy (anonymous):

89.91

OpenStudy (butterflydreamer):

yeeep :)

OpenStudy (anonymous):

Thanks!

OpenStudy (butterflydreamer):

no problem :D